# Aleth-NeRF: Illumination Adaptive NeRF with Concealing Field Assumption

### 14 Dec 2023 v1

@@ Abstract
The standard Neural Radiance Fields (NeRF) paradigm employs a viewer-centered methodology, entangling the aspects of illumination and material reflectance into emission solely from 3D points. This simplified rendering approach presents challenges in accurately modeling images captured under adverse lighting conditions, such as low light or over-exposure. Motivated by the ancient Greek emission theory that posits visual perception as a result of rays emanating from the eyes, we slightly refine the conventional NeRF framework to train NeRF under challenging light conditions and generate normal-light condition novel views unsupervisedly. We introduce the concept of a ”Concealing Field,” which assigns transmittance values to the surrounding air to account for illumination effects. In dark scenarios, we assume that object emissions maintain a standard lighting level but are attenuated as they traverse the air during the rendering process. Concealing Field thus compel NeRF to learn reasonable density and colour estimations for objects even in dimly lit situations. Similarly, the Concealing Field can mitigate over-exposed emissions during rendering stage. Furthermore, we present a comprehensive multi-view dataset captured under challenging illumination conditions for evaluation. Our code and proposed dataset is available at https://github.com/cuiziteng/Aleth-NeRF.
@@ Introduction
Can you see your days blighted by darkness ?
Pink Floyd “Lost of Words”
@@
Challenging lightness can arise from multiple sources, encompassing natural lighting variances (such as low-light situations and overly bright scenes) as well as human-induced factors (like incorrect camera exposure settings). To tackle these challenge lighting conditions, numerous techniques for image enhancement and exposure correction have been developed and proposed.
Image Enhancement &\&& Exposure Correction: Image enhancement methods aims to enhance images with poor illumination, traditional methods usually rely on RetiNex theory (Land 1986; Guo, Li, and Ling 2017) or Histogram Equalization (Gonzalez and Woods 2006), currently deep neural networks (DNNs) based methods have become the mainstream solutions, series of CNN &\&& Transformer-based methods have been developed in this direction (Wei et al. 2018; Moran et al. 2020; Wang et al. 2021b, 2022; Jiang et al. 2021; Guo et al. 2020; Jin, Yang, and Tan 2022; Ma et al. 2022; Yang et al. 2023). Meanwhile, several exposure correction methods have been proposed to consider both under &\&& over exposure conditions (Afifi et al. 2021; Nsampi, Hu, and Wang 2021; Cui et al. 2022; Huang et al. 2023), which aims to correct underexposure and its adverse overexposure images into normal-light condition. However, image enhancement &\&& exposure correction methods almost build on 2D image space operations, which often fail to exploit the 3D geometry of the scene and could not deal with multi-view inputs.
Video Enhancement &\&& Burst Enhancement: Beyond above techniques that focus on single image. Video enhancement methods have been proposed to optimize the temporal consistency between adjacent frames, ensuring stability when processing different frames. These methods employ various approaches such as 3D convolution (Lv et al. 2018), optical flow (Zhang et al. 2021a), and event guidance (Liu et al. 2023). Burst enhancement also plays a crucial role in modern computational photography area (Delbracio et al. 2021; Hasinoff et al. 2016), where multiple frames are captured during exposure and processed using an ”align-merge-enhance” approach to produce a single output frame. In recent advancements, deep neural networks have been employed to replace traditional manual operation algorithms in these methods (Godard, Matzen, and Uyttendaele 2018; Mildenhall et al. 2018; Dudhane et al. 2022).
@@ Methods
Figure 2: Train on adverse lighting condition images $\textbf{C}^{adv}$, Aleth-NeRF performs unsupervised lightness correction by (a). remove concealing fields in low-light conditions and (b). add concealing fields in over-exposure conditions.
@@
Local Concealing Field denoted as $\Omega(\textbf{r}(i))$, defines an extra light concealing capacity of a particle at 3D location $\textbf{r}(i)$. As depicted in Fig. 2, $\Omega$ is individually learned for each 3D position and is generated using the density MLP $F_{\sigma}$. To create the local concealing field $\Omega$, we introduce an additional large kernel convolution layer (with a size of 7) built upon $F_{\sigma}$. This convolution process establishes spatial relationships between pixels and enriches the Concealing Field with predominantly light-related information rather than structural information (Zamir et al. 2020). This larger kernel convolution also effectively suppresses noise and contributes to smoother rendering outcomes.
@@
Global Concealing Field denoted as $\Theta_{G}(i)$, is defined as a set of learnable parameters corresponding to the camera distance $i$ for all camera rays in R. $\Theta_{G}$ remains constant within a scene and is independent of voxels, as we posit that a specific degree of lighting influence remains consistent across the same scene. In our experiments, we initialize $\Theta_{G}(i)$ with a value of 0.5 for each camera distance $i$. With the global and local concealing fields, the accumulated transmittance $T$ in Eq. 4 is blighted by $\Omega$ and $\Theta_{G}$ to mimic the process of light suppression (see Fig. 1(b)), as follow:
$T^{conceal}(\textbf{r}(i))=\exp\left(-\sum_{j=1}^{i-1}\sigma(\textbf{r}(j))\cdot\delta\right)\cdot\prod_{j=1}^{i-1}\Omega(\textbf{r}(j))\Theta_{G}(j).$ (9)
Train and Test Schemes of Aleth-NeRF for both low-light and over-exposure scenes are illustrated in Fig. 2. For low-light conditions (blue arrow in Fig. 2), Aleth-NeRF incorporates $T^{conceal}$ during the volume rendering stage to generate low-light images $\hat{\textbf{C}}^{adv}$, while removing the concealing fields to produce normal-light images $\hat{\textbf{C}}^{nor}$ with the original $T$ in Eq. 4. Conversely, for over-exposure scenes (yellow arrow in Fig. 2), Aleth-NeRF employs $T$ to render very-bright images $\hat{\textbf{C}}^{adv}$ and incorporates the concealing fields along with $T^{conceal}$ to render normal-light images $\hat{\textbf{C}}^{nor}$. The NeRF regression loss is applied to compare predicted adverse-light images $\hat{\textbf{C}}^{adv}$ with the corresponding ground truth $\textbf{C}^{adv}$, while unsupervised lightness correction losses ensure the generation of predicted normal-light images $\hat{\textbf{C}}^{nor}$, more details please refer to next section.
Analysis of Concealing Fields is shown in Fig. 3, we analyse the distribution of concealing fields and density along the camera ray $\textbf{r}(i)$ ($z$ axis), where $x$, $y$ are the training images’ width and height. In Fig. 3, brighter region denotes the larger possibility there exists concealing fields (gray line) or density (red line). We find that concealing fields exist in location $\textbf{r}(i)$ with lower density value $\sigma(\textbf{r}(i))$. And the Pearson correlation coefficient Corr between concealing fields and density $\sigma$ are also negative along both $x-z$ axis (-0.589) and $y-z$ axis (-0.694). This validates that concealing fields are separated from density, thus rarely participating in scene rendering. Concealing fields exists more in locations $\textbf{r}(i)$ with sparse density, i.e. air outside the objects.
@@
NeRF Regression Loss: Direct apply NeRF’s MSE loss function $\mathcal{L}_{mse}$ in low-light and over-bright scenes would result in an imbalance of pixel weights. Dark pixels with small values would contribute minimally, while bright pixels with larger values would dominate the contribution. To rectify this imbalance and ensure effective NeRF training under non-standard lighting conditions, we propose the incorporation of an inverse tone curve (Brooks et al. 2019; Cui et al. 2021) denoted as $\Phi$. This curve serves to re-balance pixel weights, additionally we introduce a small value $\epsilon=1e^{-3}$ within the regression loss to facilitate more accurate estimations. As a result, the formulated inverse tone MSE loss $\mathcal{L}_{it-mse}$ for NeRF training is presented as follows:
@@
Degree &\&& Contrast Loss: In order to regulate the level of enhancement, we incorporate a degree loss denoted as $\mathcal{L}_{de}$ on the predict normal-light images $\hat{\textbf{C}}^{nor}$. $\mathcal{L}_{de}$ involves a hyperparameter $\mathbf{e}$, which signifies the enhancement degree inherent to Aleth-NeRF. To compute $\mathcal{L}_{de}$, a global average pooling operation is applied to the $\hat{\textbf{C}}^{nor}$ of each training batch. Subsequently, we minimize the degree loss $\mathcal{L}_{de}$ by comparing the pooled $\hat{\textbf{C}}^{nor}$ with the specified degree value $\mathbf{e}$. We set $\mathbf{e}$ to 0.4 in all the experiments, set $\mathbf{e}$ to other value in range of (0.35, 0.55) could also got different type desired enhancement results, ablation of $\mathbf{e}$ is in Fig. 4(b)
@@
the contrast degree loss would maintain the structure similarity between $\hat{\textbf{C}}^{nor}$ and $\textbf{C}^{adv}$, meanwhile control the contrast degree of $\hat{\textbf{C}}^{nor}$ by parameter $\eta$, ablation analyze of $\eta$ in low-light scene is shown in Fig. 4, when enhance degree $\mathbf{e}$ is fixed, higher $\eta$ represents higher contrast in $\hat{\textbf{C}}^{nor}$. While in over-exposure scene, $\mathbf{e}\cdot\eta$ would fixed to 2.
Color Constancy Loss: Photos taken in low-light &\&& over exposure conditions would easily lose their color information, directly remove &\&& add the concealing fields may cause color imbalance. To regularize the color of the predict normal-light images $\hat{\textbf{C}}^{nor}$, we introduce color constancy loss $\mathcal{L}_{cc}$ on $\hat{\textbf{C}}^{nor}$. Here we assume that $\hat{\textbf{C}}^{nor}$ obey the gray-world assumption (Buchsbaum 1980; Guo et al. 2020; Jin, Yang, and Tan 2022), as follows:
@@
Table 1: We assess novel view synthesis performance in low-light settings by comparing generated images with ground truth normal-light views. Our evaluation metrics include PSNR $\uparrow$ (indicating better quality), SSIM $\uparrow$ (indicating better structural similarity) and LPIPS $\downarrow$ (indicating closer perceptual similarity). Bold denotes the best result.
Method buu chair sofa bike shrub mean PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS NeRF (Mildenhall et al. 2020)  7.53/ 0.313/ 0.414  6.01/ 0.142/ 0.599 6.26/ 0.206/ 0.566  6.32/ 0.067/ 0.625  8.00/ 0.027/ 0.684  5.24/ 0.151/ 0.578 ① NeRF + Image Enhancement Methods NeRF + RetiNexNet (Wei et al. 2018) 16.18/ 0.754/ 0.385 16.74/ 0.731/ 0.531 15.90/ 0.826/ 0.512 17.79/ 0.654/ 0.552 15.24/ 0.283/ 0.577 16.37/ 0.657/ 0.495 NeRF + Zero-DCE (Guo et al. 2020) 14.90/ 0.678/ 0.578 15.11/ 0.615/ 0.658 17.28/ 0.713/ 0.624 14.26/ 0.547/ 0.582 12.40/ 0.384/ 0.647 14.79/ 0.587/ 0.618 NeRF + SCI (Ma et al. 2022) 13.66/ 0.769/ 0.433 14.42/ 0.736/ 0.545 18.82/ 0.813/ 0.569 12.61/ 0.560/ 0.550 16.87/ 0.411/ 0.555 15.27/ 0.658/ 0.530 NeRF + IAT (Cui et al. 2022) 14.00/ 0.652/ 0.421 14.03/ 0.706/ 0.562 10.58/ 0.532/ 0.668 16.30/ 0.660/ 0.532 10.03/ 0.149/ 0.577 12.99/ 0.540/ 0.552 ② Image Enhancement Methods + NeRF RetiNexNet (Wei et al. 2018) + NeRF 16.16/ 0.777/ 0.339 16.82/ 0.773/ 0.488 16.87/ 0.806/ 0.548 18.00/ 0.717/ 0.448 14.65/ 0.269/ 0.513 16.50/ 0.668/ 0.467 Zero-DCE (Guo et al. 2020) + NeRF 13.81/ 0.783/ 0.353 11.20/ 0.694/ 0.498 12.02/ 0.747/ 0.420 11.35/ 0.586/ 0.448 14.11/ 0.426/ 0.473 12.50/ 0.647/ 0.438 SCI (Ma et al. 2022) + NeRF 13.62/ 0.821/ 0.309 11.75/ 0.756/ 0.526 10.10/ 0.750/ 0.505 18.10/ 0.644/ 0.458 17.13/ 0.500/ 0.469 14.14/ 0.694/ 0.453 IAT (Cui et al. 2022) + NeRF 14.33/ 0.697/ 0.311 14.77/ 0.704/ 0.563 17.77/ 0.811/ 0.519 13.68/ 0.621/ 0.501 13.84/ 0.322/ 0.536 14.88/ 0.631/ 0.486 ③ Video Enhancement Methods + NeRF MBLLEN (Lv et al. 2018) + NeRF 17.34/ 0.617/ 0.493 16.24/ 0.698/ 0.502 13.69/ 0.611/ 0.488 17.41/ 0.512/ 0.550 13.17/ 0.501/ 0.555 15.57/ 0.588/ 0.518 LLVE (Zhang et al. 2021a) + NeRF 15.54/ 0.755/ 0.473 15.46/ 0.756/ 0.492 17.89/ 0.801/ 0.457 15.47/ 0.699/ 0.455 14.97/ 0.414/ 0.511 15.87/ 0.685/ 0.470 ④ Our Proposed End-to-end Method Aleth-NeRF 19.14/ 0.839/ 0.306 16.96/ 0.793/ 0.483 16.97/ 0.847/ 0.367 17.56/ 0.719/ 0.468 17.55/ 0.484/ 0.451 17.64/ 0.736/ 0.415
@@
In this section, we introduce our collected paired low-light &\&& normal-light &\&& over-exposure multi-view dataset. Previous work (Mildenhall et al. 2021) also includes some low-light scenes, but their dataset concentrates on RAW denoising rather than sRGB low-light enhancement and does not include normal-light sRGB ground truth, making it hard to evaluate enhancement performance in novel view synthesis.
We collect 5 scenes (“buu”, “chair”, “sofa”, “bike”, “shrub”) in real-world. Each scene includes $25\sim 65$ images, for each image, we generate low-light &\&& normal-light &\&& over-exposure pairs by adjusting exposure time and ISO while other configurations of the camera are fixed. We capture multi-view images by moving and rotating the camera tripod. Images are collected with resolution 3000 $\times$ 4000. We down-sample the original resolution with ratio 8 to 375 $\times$ 500 for convenience, and generate the ground truth view and angle information by adopting COLMAP (Schönberger and Frahm 2016; Schönberger et al. 2016) on the normal-light part images. For dataset split, in each scene, we choose 3 $\sim$ 5 images as the testing set, 1 image as the validation set, and other images to be the training set. More details of our dataset please refer to supplementary.

### 24 Jan 2024 v3

@@ Abstract
The standard Neural Radiance Fields (NeRF) paradigm employs a viewer-centered methodology, entangling the aspects of illumination and material reflectance into emission solely from 3D points. This simplified rendering approach presents challenges in accurately modeling images captured under adverse lighting conditions, such as low light or over-exposure. Motivated by the ancient Greek emission theory that posits visual perception as a result of rays emanating from the eyes, we slightly refine the conventional NeRF framework to train NeRF under challenging light conditions and generate normal-light condition novel views unsupervisedly. We introduce the concept of a “Concealing Field,” which assigns transmittance values to the surrounding air to account for illumination effects. In dark scenarios, we assume that object emissions maintain a standard lighting level but are attenuated as they traverse the air during the rendering process. Concealing Field thus compel NeRF to learn reasonable density and colour estimations for objects even in dimly lit situations. Similarly, the Concealing Field can mitigate over-exposed emissions during rendering stage. Furthermore, we present a comprehensive multi-view dataset captured under challenging illumination conditions for evaluation. Our code and proposed dataset are available at https://github.com/cuiziteng/Aleth-NeRF.
@@ Introduction
Challenging lightness can arise from multiple sources, encompassing natural lighting variances (such as low-light situations and overly bright scenes) as well as human-induced factors (such as incorrect camera exposure settings). To tackle these challenge lighting conditions, numerous techniques for image enhancement and exposure correction have been developed and proposed.
Image Enhancement &\&& Exposure Correction:
Image enhancement methods aims to enhance images with poor illumination, traditional methods usually rely on RetiNex theory (Land 1986; Guo, Li, and Ling 2017) or Histogram Equalization (Gonzalez and Woods 2006), currently deep neural networks (DNNs) based methods become the mainstream solutions, series of CNN &\&& Transformer-based methods have been developed (Wei et al. 2018; Moran et al. 2020; Wang et al. 2021b, 2022; Jiang et al. 2021; Guo et al. 2020; Jin, Yang, and Tan 2022; Ma et al. 2022; Yang et al. 2023). Meanwhile, several exposure correction methods have been proposed to consider both under &\&& over exposure conditions (Afifi et al. 2021; Nsampi, Hu, and Wang 2021; Cui et al. 2022; Huang et al. 2023), which aims to correct underexposure and its adverse overexposure images into normal-light condition. However, image enhancement &\&& exposure correction methods almost build on 2D image space operations, which often fail to exploit the 3D geometry of the scene and could not deal with multi-view inputs.
Video Enhancement &\&& Burst Enhancement:
Beyond above techniques that focus on single image. Video enhancement methods have been proposed to optimize the temporal consistency between adjacent frames, ensuring stability when processing different frames. These methods employ various approaches such as 3D convolution (Lv et al. 2018), optical flow (Zhang et al. 2021a), and event guidance (Liu et al. 2023). Burst enhancement also plays a crucial role in modern computational photography area (Delbracio et al. 2021; Hasinoff et al. 2016), where multiple frames are captured during exposure and processed using an ”align-merge-enhance” approach to produce a single output frame. In recent advancements, deep neural networks have been employed to replace traditional manual operation algorithms in these methods (Godard, Matzen, and Uyttendaele 2018; Mildenhall et al. 2018; Dudhane et al. 2022).
@@ Methods
Figure 2: Train on adverse lighting condition images $C^{adv}$, Aleth-NeRF performs unsupervised lightness correction by (a). remove concealing fields in low-light conditions and (b). add concealing fields in over-exposure conditions.
@@
Local Concealing Field
denoted as $\Omega(\textbf{r}(i))$, defines an extra light concealing capacity of a particle at 3D location $\textbf{r}(i)$. As depicted in Fig. 2, $\Omega$ is individually learned for each 3D position and is generated using the density MLP $F_{\sigma}$. To create the local concealing field $\Omega$, we introduce an additional large kernel convolution layer (with a size of 7) built upon $F_{\sigma}$. This convolution process establishes spatial relationships between pixels and enriches the Concealing Field with predominantly light-related information rather than structural information (Zamir et al. 2020). This larger kernel convolution also effectively suppresses noise and contributes to smoother rendering outcomes.
@@
Global Concealing Field
denoted as $\Theta_{G}(i)$, is defined as a set of learnable parameters corresponding to the camera distance $i$ for all camera rays in R. $\Theta_{G}$ remains constant within a scene and is independent of voxels, as we posit that a specific degree of lighting influence remains consistent across the same scene. In our experiments, we initialize $\Theta_{G}(i)$ with a value of 0.5 for each camera distance $i$. With the global and local concealing fields, the accumulated transmittance $T$ in Eq. 4 is blighted by $\Omega$ and $\Theta_{G}$ to mimic the process of light suppression (see Fig. 1(b)), as follow:
$T^{conceal}(\textbf{r}(i))=$ $\exp\left(-\sum_{j=1}^{i-1}\sigma(\textbf{r}(j))\cdot\delta\right)$ (9)
$\cdot\prod_{j=1}^{i-1}\Omega(\textbf{r}(j))\Theta_{G}(j).$
Train and Test Schemes
of Aleth-NeRF for both low-light and over-exposure scenes are illustrated in Fig. 2. For low-light conditions (blue arrow in Fig. 2), Aleth-NeRF incorporates $T^{conceal}$ during the volume rendering stage to generate low-light images $\hat{\textbf{C}}^{adv}$, while removing the concealing fields to produce normal-light images $\hat{\textbf{C}}^{nor}$ with the original $T$ in Eq. 4. Conversely, for over-exposure scenes (yellow arrow in Fig. 2), Aleth-NeRF employs $T$ to render very-bright images $\hat{\textbf{C}}^{adv}$ and incorporates the concealing fields along with $T^{conceal}$ to render normal-light images $\hat{\textbf{C}}^{nor}$. The NeRF regression loss is applied to compare predicted adverse-light images $\hat{\textbf{C}}^{adv}$ with the corresponding ground truth $\textbf{C}^{adv}$, while unsupervised lightness correction losses ensure the generation of predicted normal-light images $\hat{\textbf{C}}^{nor}$, more details please refer to next section.
Analysis of Concealing Fields
is shown in Fig. 3, we analyse the distribution of concealing fields and density along the camera ray $\textbf{r}(i)$ ($z$ axis), where $x$, $y$ are the training images’ width and height. In Fig. 3, brighter region denotes the larger possibility there exists concealing fields (gray line) or density (red line). We find that concealing fields exist in location $\textbf{r}(i)$ with lower density value $\sigma(\textbf{r}(i))$. And the Pearson correlation coefficient Corr between concealing fields and density $\sigma$ are also negative along both $x-z$ axis (-0.589) and $y-z$ axis (-0.694). This validates that concealing fields are separated from density, thus rarely participating in scene rendering. Concealing fields exists more in locations $\textbf{r}(i)$ with sparse density, i.e. air outside the objects.
@@
NeRF Regression Loss:
Direct apply NeRF’s MSE loss function $\mathcal{L}_{mse}$ in low-light and over-bright scenes would result in an imbalance of pixel weights. Dark pixels with small values would contribute minimally, while bright pixels with larger values would dominate the contribution. To rectify this imbalance and ensure effective NeRF training under non-standard lighting conditions, we propose the incorporation of an inverse tone curve (Brooks et al. 2019; Cui et al. 2021) denoted as $\Phi$. This curve serves to re-balance pixel weights, additionally we introduce a small value $\epsilon=1e^{-3}$ within the regression loss to facilitate more accurate estimations. As a result, the formulated inverse tone MSE loss $\mathcal{L}_{it-mse}$ for NeRF training is presented as follows:
@@
Degree &\&& Contrast Loss:
In order to regulate the level of enhancement, we incorporate a degree loss denoted as $\mathcal{L}_{de}$ on the predict normal-light images $\hat{\textbf{C}}^{nor}$. $\mathcal{L}_{de}$ involves a hyperparameter $\mathbf{e}$, which signifies the enhancement degree inherent to Aleth-NeRF. To compute $\mathcal{L}_{de}$, a global average pooling operation is applied to the $\hat{\textbf{C}}^{nor}$ of each training batch. Subsequently, we minimize the degree loss $\mathcal{L}_{de}$ by comparing the pooled $\hat{\textbf{C}}^{nor}$ with the specified degree value $\mathbf{e}$. We set $\mathbf{e}$ to 0.45 in all experiments, ablation analyse of $\mathbf{e}$ is in Fig. 4(b).
@@
the contrast degree loss would maintain the structure similarity between $\hat{\textbf{C}}^{nor}$ and $\textbf{C}^{adv}$, meanwhile control the contrast degree of $\hat{\textbf{C}}^{nor}$ by parameter $\eta$, ablation analyze of $\eta$ in low-light scene is shown in Fig. 4, when enhance degree $\mathbf{e}$ is fixed, higher $\eta$ represents higher contrast in $\hat{\textbf{C}}^{nor}$. While in over-exposure scene, $\mathbf{e}\cdot\eta$ would fixed to $0.5$.
Color Constancy Loss:
Photos taken in low-light &\&& over exposure conditions would easily lose their color information, directly remove &\&& add the concealing fields may cause color imbalance. To regularize the color of the predict normal-light images $\hat{\textbf{C}}^{nor}$, we introduce color constancy loss $\mathcal{L}_{cc}$ on $\hat{\textbf{C}}^{nor}$. Here we assume that $\hat{\textbf{C}}^{nor}$ obey the gray-world assumption (Buchsbaum 1980; Guo et al. 2020; Jin, Yang, and Tan 2022), as follows:
@@
In this section, we introduce our collected paired low-light &\&& normal-light &\&& over-exposure multi-view dataset, names LOM dataset. Previous work (Mildenhall et al. 2021) also includes some low-light scenes, but their dataset concentrates on RAW denoising rather than sRGB low-light enhancement and does not include normal-light sRGB ground truth, making it hard to evaluate enhancement performance in novel view synthesis.
scene buu chair sofa bike shrub collected views 25 48 33 40 35 training views 22 43 29 36 30 evaluation views 3 5 4 4 5
Table 1: Details of the dataset split for LOM.
In our proposed LOM dataset, we collected 5 scenes (“buu”, “chair”, “sofa”, “bike”, “shrub”) in real-world. Each scene includes $25\sim 48$ images, for each view’s image, we generate low-light &\&& normal-light &\&& over-exposure pairs by adjusting exposure time and ISO while other configurations of the camera are fixed. We capture multi-view images by moving and rotating the camera tripod. Images are collected with resolution 3000 $\times$ 4000. We down-sample the original resolution with ratio 8 to 375 $\times$ 500 for convenience, and generate the ground truth view and angle information by adopting COLMAP (Schönberger and Frahm 2016; Schönberger et al. 2016) on the normal-light part images. For dataset split, in each scene, we choose 3 $\sim$ 5 images as the testing set, 1 image as the validation set, and other images to be the training set, details of training and evaluation views split is shown in Table. 1. More details of our dataset please refer to supplementary part.
@@
Method buu chair sofa bike shrub mean PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS NeRF  7.51/ 0.291/ 0.448  6.04/ 0.147/ 0.594 6.28/ 0.210/ 0.568  6.35/ 0.072/ 0.623  8.03/ 0.031/ 0.680  6.84/ 0.150/ 0.582 ① NeRF + Image Enhancement Methods NeRF + RetiNexNet 11.64/ 0.646/ 0.395 11.04/ 0.631/ 0.561 12.85/ 0.738/ 0.527 17.79/ 0.653/ 0.550 11.85/ 0.211/ 0.583 13.03/ 0.576/ 0.523 NeRF + Zero-DCE 17.81/ 0.833/ 0.357 12.44/ 0.684/ 0.547 14.43/ 0.787/ 0.539 10.16/ 0.468/ 0.557 12.58/ 0.282/ 0.540 13.48/ 0.610/ 0.488 NeRF + SCI 7.84/ 0.660/ 0.562 12.07/ 0.699/ 0.584 10.25/ 0.737/ 0.626 18.84/ 0.637/ 0.565 12.38/ 0.358/ 0.587 12.27/ 0.618/ 0.585 NeRF + IAT 14.03/ 0.656/ 0.453 19.08/ 0.800/ 0.565 10.49/ 0.528/ 0.678 17.16/ 0.657/ 0.534 16.15/ 0.344/ 0.573 15.38/ 0.597/ 0.561 ② Image Enhancement Methods + NeRF RetiNexNet + NeRF 16.19/ 0.780/ 0.396 16.89/ 0.756/ 0.543 16.98/ 0.807/ 0.577 18.00/ 0.707/ 0.482 14.86/ 0.284/ 0.518 16.58/ 0.667/ 0.503 Zero-DCE + NeRF 17.90/ 0.858/ 0.376 12.58/ 0.721/ 0.460 14.45/ 0.831/ 0.419 10.39/ 0.518/ 0.464 12.32/ 0.308/ 0.481 13.53/ 0.649/ 0.432 SCI + NeRF 7.76/ 0.692/ 0.525 19.77/ 0.802/ 0.674 10.08/ 0.772/ 0.520 13.44/ 0.658/ 0.435 18.16/ 0.503/ 0.475 13.84/ 0.689/ 0.510 IAT + NeRF 14.46/ 0.705/ 0.386 18.70/ 0.780/ 0.665 17.88/ 0.829/ 0.547 13.65/ 0.616/ 0.528 13.87/ 0.317/ 0.536 15.71/ 0.649/ 0.532 ③ Video Enhancement Methods + NeRF MBLLEN + NeRF 22.39/ 0.877/ 0.353 23.59/ 0.788/ 0.559 19.99/ 0.836/ 0.542 14.09/ 0.636/ 0.525 13.17/ 0.501/ 0.555 18.65/ 0.728/ 0.507 LLVE + NeRF 19.97/ 0.848/ 0.393 15.17/ 0.764/ 0.610 18.17/ 0.855/ 0.465 13.84/ 0.638/ 0.492 15.35/ 0.287/ 0.577 16.50/ 0.678/ 0.507 ④ Our Proposed End-to-end Method Aleth-NeRF 20.22/ 0.859/ 0.315 20.93/ 0.818/ 0.468 19.52/ 0.857/ 0.354 20.46/ 0.727/ 0.499 18.24/ 0.511/ 0.448 19.87/ 0.754/ 0.417
Table 2: We assess novel view synthesis performance in low-light settings by comparing generated images with ground truth normal-light views. Our evaluation metrics include PSNR $\uparrow$, SSIM $\uparrow$ and LPIPS $\downarrow$. Bold denotes the best result, Underline denotes the second best result.
Method buu chair sofa bike shrub mean PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS NeRF  7.12/0.674/0.499 11.05/ 0.741/ 0.418 10.22/ 0.783/ 0.475  9.65/ 0.698/ 0.416  9.96/ 0.405/ 0.480  9.60/ 0.660/ 0.457 ① NeRF + Exposure Correction Methods NeRF + HE 14.34/ 0.613/ 0.673 15.37/ 0.661/ 0.590 16.69/ 0.733/ 0.558 15.32/ 0.627/ 0.458 11.97/ 0.468/ 0.556 14.74/ 0.620/ 0.567 NeRF + IAT 14.11/ 0.780/ 0.433 19.24/ 0.810/ 0.491 16.60/ 0.837/ 0.459 17.73/ 0.760/ 0.394 14.05/ 0.381/ 0.499 16.35/ 0.714/ 0.455 NeRF + MSEC 16.13/ 0.800/ 0.427 15.60/ 0.786/ 0.472 16.56/ 0.807/ 0.495 12.60/ 0.716/ 0.465 13.66/ 0.332/ 0.509 14.91/ 0.688/ 0.473 ② Exposure Correction Methods + NeRF HE + NeRF 14.65/ 0.743/ 0.519 15.55/ 0.736/ 0.497 17.11/ 0.781/ 0.477 15.77/ 0.692/ 0.367 12.61/ 0.506/ 0.622 15.13/ 0.692/ 0.496 IAT + NeRF 16.22/ 0.815/ 0.486 18.98/ 0.799/ 0.503 18.45/ 0.849/ 0.478 19.63/ 0.776/ 0.408 15.63/ 0.434/ 0.477 17.78/ 0.734/ 0.470 MSEC + NeRF 15.53/ 0.817/ 0.499 16.95/ 0.758/ 0.580 19.60/ 0.817/ 0.498 18.90/ 0.725/ 0.483 15.48/ 0.400/ 0.499 17.29/ 0.703/ 0.512 ③ Our Proposed End-to-end Method Aleth-NeRF 16.78/ 0.805/ 0.611 20.08/ 0.820/ 0.499 17.85/ 0.852/ 0.458 19.85/ 0.773/ 0.392 15.91/ 0.477/ 0.483 18.09/ 0.745/ 0.488
Table 3: We assess novel view synthesis performance in over-exposure settings by comparing generated images with ground truth normal-light views. Our evaluation metrics include PSNR $\uparrow$, SSIM $\uparrow$ and LPIPS $\downarrow$. Bold denotes the best result, Underline denotes the second best result.